# Topological states constructed by two different trivial quantum wires

### v1

@@ III Three-leg ladder
Figure 4: (a)The lattice sites of the type I three-leg ladder is marked as $A$, $B$, $C$, and $D$. (b)The lattice sites of the type II three-leg ladder is marked as $A$, $B$, $C$, $D$, and $E$.
@@
Motivated and encouraged by the above results, we investigate the three-leg ladder system. When the two-leg ladder is coupled with the third trivial quantum wire, we could obtain two different types of three-leg ladders. In the type I three-leg ladder (Fig.3a), the lattice constant of center chain is twice that of top chain and bottom chain. In the type II three-leg ladder (Fig.3b), the lattice constant of top chain and bottom chain are twice that of center chain.
@@
Thus, the whole Hilbert space becomes separable. It means that the block diagonalization separates the Hilbert space of the three-leg ladder into two independently subspaces. The $h_{1}$ is a trivial subspace. The $h_{2}$ is a topological nontrivial subspace, which is similar to the Hilbert space of the asymmetric two-leg model (Eq.15) with the effective inter-chain hoppings $\sqrt{2}t_{1}$ and $\sqrt{2}t_{2}$. Thus, the analysis method of section II.2 can be applied to this nontrivial subspace.
@@
where the unitary matrix $Q_{2}$ and $h_{3}$
@@
Next, we numerically confirm the existence of topological insulators and topological metals and the topological edge states in the two three-leg ladders. In Fig.5a and Fig.5e, we plot the energy spectrum of the two three-leg ladders under OBC as a function of phase factor $\phi$, when the system parameters are $L=60,J=1,J_{1}=1,\lambda=6$. The Chern numbers of $h_{2}$ for the three bands are $(C_{1},C_{3},C_{3})=(1,-1,0)$. The nonzero Chern bands will induce different topological states in the two three-leg ladders. For the type I three-leg ladder, we plot the energy spectrum in the real-space under OBC when the system parameters are $L=60,J_{1}=1,J_{2}=1+6\cos(3\pi/4),t_{1}=1+6\cos(17\pi/12),t_{2}=1+6\cos(25\pi/12)$, and $\phi=3\pi/4$. It is easy to see that there are two topological edge states (red circles) with different energies appear. At filling $1/3$, the three-leg ladder is in topological insulator phase with one topological edge state. As the filling factor increases, the system becomes a topological metal state (zero energy around). The spatial distribution of the two topological edge states is shown in Fig.5(c,d). For the type II three-leg ladder, the energy spectrum of the type II three-leg ladder in the real-space under OBC is shown in Fig.5f, when the system parameters are $L=60,J_{1}=1,J_{2}=7,t_{1}=1+6\cos(2\pi/3),t_{2}=1+6\cos(4\pi/3)$, and $\phi=0$. In this case, inversion symmetry is recovered and the system becomes a topological metal state. At the same time, the two degenerate edge states (red circles) appear, whose density profiles are illustrated in Fig.5(g,h).
IV Summary and outlook
@@
In short, we use the two-leg and three-leg ladders formed by the two different trivial quantum wires to construct the topological states. For the two-leg ladder, the system can emerge the inversion symmetry topological insulator with two degenerate topological edge states, the inversion-broken topological insulator with one or two topological edge state of different energies, and topological metals with edge states embedded in the bulk states (topological bound states in the continuum). For the three-leg ladder, using a unitary transformation, we separate the Hilbert space of the system into a trivial subspace and a topological nontrivial subspace. The nontrivial subspace corresponds to the whole Hilbert space of the asymmetric two-leg ladder. The topological insulators and topological metals dependent on the filling factor can also appear in the three-leg ladder. On the other hand, these topological states for the two-leg and three-leg ladders could be realized experimentally by a photonic waveguide array, photonic crystals, topoelectrical circuits, or coupled acoustic resonators. In addition, based on the results of this work, we believe that the constructions for two-dimensional and three-dimensional quantum systems using two or more different trivial quantum wires may realize rich topological states.

### v2

@@ III Three-leg ladder
Figure 5: (a)The lattice sites of the type I three-leg ladder is marked as $A$, $B$, $C$, and $D$. (b)The lattice sites of the type II three-leg ladder are marked as $A$, $B$, $C$, $D$, and $E$.
@@
Motivated and encouraged by the above results, we investigate the three-leg ladder system. When the two-leg ladder is coupled with the third trivial quantum wire, we can obtain two different types of three-leg ladders. In the type I three-leg ladder (Fig.3a), the lattice constant of center chain is twice that of top chain and bottom chain. In the type II three-leg ladder (Fig.3b), the lattice constant of top chain and bottom chain are twice that of center chain.
@@
Thus, the whole Hilbert space becomes separable. It means that the block diagonalization separates the Hilbert space of the three-leg ladder into two independent subspaces. The $h_{1}$ is a trivial subspace. The $h_{2}$ is a topological nontrivial subspace, which is similar to the Hilbert space of the asymmetric two-leg model (Eq.6) with the effective inter-chain hoppings $\sqrt{2}t_{1}$ and $\sqrt{2}t_{2}$. Thus, the analysis method of section II.2 can be applied to this nontrivial subspace.
@@
where the unitary matrix $Q_{2}$ and $h_{3}$ are
@@
Next, we numerically confirm the existence of topological insulators and topological metals and the topological edge states in the two three-leg ladders. In Fig.6a and Fig.6e, we plot the energy spectrum of the two three-leg ladders under OBC as a function of phase factor $\phi$, when the system parameters are $L=60,J=1,J_{1}=1,\lambda=6$. The Chern numbers of $h_{2}$ for the three bands are $(C_{1},C_{2},C_{3})=(1,-1,0)$. The nonzero Chern bands will induce different topological states in the two three-leg ladders. For the type I three-leg ladder, we plot the energy spectrum in the real-space under OBC when the system parameters are $L=60,J_{1}=1,J_{2}=1+6\cos(3\pi/4),t_{1}=1+6\cos(17\pi/12),t_{2}=1+6\cos(25\pi/12)$, and $\phi=3\pi/4$. It is easy to see that there are two topological edge states (red circles) with different energies appear. At filling $1/3$, the three-leg ladder is in topological insulator phase with one topological edge state. As the filling factor increases, the system becomes a topological metal state (zero energy around). The spatial distribution of the two topological edge states is shown in Fig.6(c,d). For the type II three-leg ladder, the energy spectrum of the type II three-leg ladder in the real-space under OBC is shown in Fig.6f, when the system parameters are $L=60,J_{1}=1,J_{2}=7,t_{1}=1+6\cos(2\pi/3),t_{2}=1+6\cos(4\pi/3)$, and $\phi=0$. In this case, inversion symmetry is recovered and the system becomes a topological metal state. At the same time, the two degenerate edge states (red circles) appear, whose density profiles are illustrated in Fig.6(g,h).
IV Two-leg ladder formed by quantum wires with other lattice constant ratios
Figure 7: (a) The lattice sites of the two-leg ladder are labeled by $A$, $B$, $C$, and $D$. The strength of intra-chain interactions are $J_{1}$ and $J_{2}$, and the inter-chain coupling amplitudes are $t_{1}$ and $t_{2}$.(b) Energy spectra as a function of inter-chain hopping amplitude $t_{1}$ with system size $L=120$, $J_{1}=t_{2}=1$, and $J_{2}=2$ under OBC. The red dots indicate the inversion-symmetry-protected topological edge modes. (c) The density profiles of the two degenerate topological bound edge states for filling $1/2$, other parameters are $L=120$, $J_{1}=t_{2}=1$, $J_{2}=2$, and $t_{1}=0.5$.
Now, we investigate the symmetric intra-chain hopping two-leg ladder formed by two trivial quantum wires with other lattice constant ratios. To begin with, we analyze a two-leg ladder case, where the lattice constant of the upper chain is three times that of the bottom chain, which is illustrated in Fig.7a. The Hamiltonian of the tight-binding model in real space could be written as:
$H_{a}=\!\!\!\!\!\!\!\!\!$ $\sum_{j}^{L}(t_{1}c^{\dagger}_{A,j}c_{B,j}+t_{2}c^{\dagger}_{A,j}c_{C,j}+t_{1}c^{\dagger}_{A,j}c_{D,j}$ (27)
$+J_{2}c^{\dagger}_{B,j}c_{C,j}+J_{2}c^{\dagger}_{C,j}c_{D,j})$
$+\sum_{j}^{L-1}(J_{1}c^{\dagger}_{A,j}c_{A,j+1}+J_{2}c^{\dagger}_{D,j}c_{B,j+1})+h.c.\ ,$
Then, we can express bulk Hamiltonian matrix in momentum space as (base $\psi^{\dagger}_{k}=(c^{\dagger}_{k,A},c^{\dagger}_{k,B},c^{\dagger}_{k,C},c^{\dagger}_{k,D})$):
$H_{a}(k)=\left[\begin{matrix}2J_{1}\cos(k)&t_{1}&t_{2}&t_{1}\\ t_{1}&0&J_{2}&J_{2}e^{ik}\\ t_{2}&J_{2}&0&J_{2}\\ t_{1}&J_{2}e^{-ik}&J_{2}&0\\ \end{matrix}\right].$ (28)
For this system, the inversion symmetry operator $\mathcal{I}$ is defined by
$\mathcal{I}_{a}=\left[\begin{matrix}1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ \end{matrix}\right].$ (29)
We present the OBC energy spectrum of the finite system in Fig.7b, for $J_{1}=t_{2}=1$ and $J_{2}=2$. When the filling factor is $1/2$, the nontrivial topological invariant $\mathcal{N}=1$ indicates that a pair of topologically protected degenerate in-gap edge modes will appear at the system boundaries. The topologically localized boundary modes of nontrivial topological insulators are indicated by red dots. The bulk spectrum is gapped for filling $1/2$. As shown in Fig.7c, the two degenerate nontrivial edge states with $L=120$, $J_{1}=t_{2}=1$, $J_{2}=2$, and $t=0.5$ are clearly localized in the left and right boundaries.
Figure 8: (a) The lattice sites of the unit cell for the two-leg ladder are termed $A$, $B$, $C$, $D$, and $E$. The strength of intra-chain interactions are $J_{1}$ and $J_{2}$, and the inter-chain coupling amplitudes are $t_{1}$ and $t_{2}$. The phase diagram of the two-leg ladder with $J_{1}=1$ and $J_{2}=2$ at bands as (b) lowest energy band, (c) two lower energy bands, (d) three lower energy bands, and (e) four lower energy bands. The yellow (red) region indicates that the system is in the nontrivial (trivial) topological insulator phase or topological metal phase. (f) The energy spectra under OBC as a function of inter-chain coupling amplitude $t_{1}$ with system size $L=60$, $J_{1}=t_{2}=1$, and $J_{2}=2$. The red dots indicate the topologically protected edge modes. (g) Bloch band structures ($t_{1}=0.5$) in subgraph (f) indicate the indirect bandgap closure between the second and third bands. (h,i) The density profiles of the topological bound edge states of topological metal, other parameters are $L=60$, $J_{1}=t_{2}=1$, $J_{2}=2$, and $t_{1}=0.5$.
Next, we investigate another two-leg ladder system (illustrated in Fig.8a) in that the lattice constant of the upper chain is one and a half times that of the bottom chain. The tight-binding Hamiltonian in real space can be described by
$H_{b}=\sum_{j}^{L}(t_{1}c^{\dagger}_{A,j}c_{B,j}+t_{2}c^{\dagger}_{A,j}c_{C,j}+t_{1}c^{\dagger}_{D,j}c_{E,j}$
$+t_{2}c^{\dagger}_{C,j}c_{E,j}+J_{1}c^{\dagger}_{A,j}c_{E,j}+J_{2}c^{\dagger}_{B,j}c_{C,j}+J_{2}c^{\dagger}_{C,j}c_{D,j})$
$+\sum_{j}^{L-1}(J_{1}c^{\dagger}_{E,j}c_{A,j+1}+J_{2}c^{\dagger}_{D,j}c_{B,j+1})+h.c.$ (30)
where $c^{\dagger}_{\alpha,j}(c_{\alpha,j})$ denotes the creation (annihilation) operator at site $\alpha$ ($\alpha$ stands for the lattice sites $A$, $B$, $C$, $D$, or $E$.) of the $j$-th unit cell.
Then, the bulk Hamiltonian matrix in momentum space (base $\psi^{\dagger}_{k}=(c^{\dagger}_{k,A},c^{\dagger}_{k,B},c^{\dagger}_{k,C},c^{\dagger}_{k,D},c^{\dagger}_{k,E})$) could be written
$\!\!\!\!\!\!\!\!\!H_{b}(k)\!=\!\!\left[\begin{matrix}0&t_{1}&t_{2}&0&J_{1}+J_{1}e^{ik}\\ t_{1}&0&J_{2}&J_{2}e^{ik}&0\\ t_{2}&J_{2}&0&J_{2}&t_{2}\\ 0&J_{2}e^{-ik}&J_{2}&0&t_{1}\\ J_{1}+J_{1}e^{-ik}&0&t_{2}&t_{1}&0\\ \end{matrix}\right]\!.$ (31)
The existence of edge states in this tight-binding model can be explained by nontrivial inversion symmetric topological invariant $\mathcal{N}$. The phase diagram for the invariant $\mathcal{N}$ of the two-leg ladder in the $t_{1}-t_{2}$ parameter space for distinct bands: (a) lowest energy band, (b) two lower energy bands, (c) three lower energy bands, and (d) four lower energy bands, which are shown in Fig.8 b-e, respectively. The yellow region ($\mathcal{N}=1$, the Zak phase is $\pi$) represents the existence of topological insulator phase or topological metal phase in the system. The red region ($\mathcal{N}=0$, the Zak phase is 0) represents the existence of trivial insulator or metal in the model. In topological insulator phase, a pair of topologically protected degenerate edge states would appear in the finite system under OBC. The energy spectrum of the finite system under OBC is presented in Fig.8f, for $J_{1}=t_{2}=1$ and $J_{2}=2$. For small inter-chain hopping $t_{1}$, the second band and third band overlap, the system becomes inversion symmetry protected topological metal when the Fermi energy fixed between these two bands. In Fig.8g, the bulk band structures ($t_{1}=0.5$) indicate the indirect nature of bandgap closure between the second and third bands. As shown in Fig.8(h,i), the nontrivial edge states of inversion symmetry protected topological metal with $L=60$, $J_{1}=t_{2}=1$, $J_{2}=2$, and $t_{1}=0.5$ are clearly localized in the boundaries. As the hopping amplitude $t_{1}$ increases, when the bulk spectrum is gapped for filling factor $2/5$, the system changes from the inversion symmetric topological insulator with nontrivial edge states to trivial insulator. At fillings $1/5$, and $4/5$ the bulk spectrum are gapped. As the hopping amplitude $t_{1}$ increases, the system changes from trivial insulator to the inversion symmetric topological insulator with nontrivial edge states. However, for filling $3/5$, the system evolves from inversion symmetric topological insulator to trivial insulator, as the hopping $t_{1}$ increases.
When the inversion symmetry is broken in the two two-leg ladders in Figs.7 and 8, we can use the same method (mapping the systems to AAH chains) of subsection II.2 to analyze the topological properties. It is straightforward to analyze and inspect the rich topological phases such as topological insulators and topological metals.
V Summary and outlook
@@
In short, we use the two-leg and three-leg ladders formed by the two different trivial quantum wires to construct the topological states. For the two-leg ladder formed by two different trivial quantum wires with a lattice constant ratio of 1:2, the system can emerge the inversion symmetric topological insulator with two degenerate topological edge states, the inversion-broken topological insulator with one or two topological edge states of different energies, and topological metals with edge states embedded in the bulk states (topological bound states in the continuum). For the three-leg ladder, using a unitary transformation, we separate the Hilbert space of the system into a trivial subspace and a topological nontrivial subspace. The nontrivial subspace corresponds to the whole Hilbert space of the asymmetric two-leg ladder. The topological insulators and topological metals dependent on the filling factor can also appear in the three-leg ladder. The constructions for a two-leg ladder using two different trivial quantum wires with a lattice constant ratio of 1:3 and 2:3 could realize the inversion symmetric topological insulator and topological metals with topological edge states. On the other hand, these topological states for the two-leg and three-leg ladders could be realized experimentally by a photonic waveguide array, photonic crystals, topoelectrical circuits, or coupled acoustic resonators. In addition, based on the results of this work, we believe that the constructions for 2D and 3D quantum systems using two or more different trivial quantum wires may realize rich topological states. Furthermore, constructing a two-leg ladder using two trivial quantum wires with a commensurate lattice constant ratio may realize 1D moiré superlattice structures[76]. Lastly, using a quasiperiodic incommensurate lattice constant ratio may create 1D bichromatic incommensurate lattices[77, 78]. These constructed structures could emerge elegant physical features.